OpenStudy (anonymous):

((6x^4)/(4x^9))((12x^5)/3x^2)) what is the answer to this? A) 6/x^2 B) 6x^8 C) 6/x^8 D) 6x^2

OpenStudy (anonymous):

x^a * x^b = x^(a+b) x^a / x^b = x^(a-b)

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

these are the two rules you need...multiply all the terms in the denominator and numerator separately first

OpenStudy (anonymous):

i thought the answer was c

OpenStudy (anonymous):

am i right?